Statistics Expressing the Perception of Pastors on Mission
Strongly Agree = 1, Agree = 2, Neutral = 3, Disagree = 4, Strongly Disagree = 5
Item/Statement Mean
Mode Median Standard Deviation
(SD)
1. Every RPAG single member has the responsibility of fulfilling the Great Commission and must participate in funding this effort aside of paying regular tithes and offerings. 1.31 1 1 0.54
2. Jesus called RPAG churches to take the Good News to the whole world. 1.39 1 1 0.67
6. As a pastor, one of my great responsibilities is to prepare my church for the work of mission in foreign lands. 1.53 1 1 0.81
Cluster Mean, and SD
1.41 0.67
Interpret the mean and standard deviation
1
Expert's answer
2020-06-30T11:46:01-0400

For question all the questions pastors 95% of pastors agree with all the 3 statements since we know from the empirical rule that 95% = mean ± 2standard deviations

so calculations will be 95% of data falls between 1.31 ± 2*0.54 = (0.23, 2.4)

For second question 95% of data falls between 1.39 ± 2*0.67 = (0.05, 2.73)

For third question 95% of data falls between 1.53 ± 2*0.81 = (-0.09, 3.15)

Average for all questions 95% of data falls between 1.41 ± 2*0.67 = (0.07, 2.75)

From the above calculations we can conclude that 95% of pastors agree with the above statements.

we can also say pastors differed in their perception in statement compared to statement one and two.